OpenStudy (anonymous):

several friends charted a boat for the days fishing. They caught a total of 68 fish. Determine the empirical probability the next fish caught is a King Fish. Table Shows 16 Grouper, 17 Sharks, 2 flounders and 33 King Fish

OpenStudy (tkhunny):

Have you considered counting all the fish and calculating the empirical probabilities?

OpenStudy (tkhunny):

Feel free to answer on this thread. Empirical Probabilities are just those that are observed. They do not require theories. How many fish in total?

OpenStudy (anonymous):

68

OpenStudy (tkhunny):

Perfect. 2/68 = 0.0294 -- This is the empirical probability of catching a flounder, given that you caught something. You do the King Fish, since that's what the problem statement wants.

OpenStudy (anonymous):

0.485...

OpenStudy (tkhunny):

That's what I get. That's the arithmetic. Can you now answer the question?

OpenStudy (anonymous):

so I just divide the number of fish by the total to get the awnser

OpenStudy (tkhunny):

Please don't do any more dividing. You already calculated the right values. 0.485... is the empirical probability of catching a King Fish, given that you caught something. What does the problem statement ask?

OpenStudy (anonymous):

no I was saying..when dealing with empirical probability do I always divide the sum into the total

OpenStudy (tkhunny):

That is an excellent question. The point of a Probability Distribution is to cover ALL possibilities - 100% of the probability. Check it out. Calculate the individual empirical probabilities. Grouper 16/68 = 0.23529 Sharks 17/68 = 0.25000 Flounders 2/68 = 0.02941 King Fish 33/68 = 0.48530 0.23529 + 0.25000 + 0.02941 + 0.48530 = 1.00000 -- Pretty cool?